5.

\textbf{F}(x,y,z) = y\textbf{j} + z\textbf{k}

Let E be the solid bounded by the paraboloid y = x2 + z2 , the cylinder x2 + z2 = 1, and the plane y = 2. Let S be the surface of E with outward orientation.

(b) Evaluate the volume integral

\iiint_{E}^{ } \triangledown \cdot \textbf{F}dV.
Answer #1

F(xcY, 2) yie2k VF = 60 4 By + 2 k) (yj 12 k) DIF = (1+1)-2 given cuaves y = x + 22, 2+2 2 , and phone y = converting to cyli